A STUDY OF THE FO RADICAL BY FAR-IR LMR

\underline{FILIPPO TAMASSIA}, JOHN M. BROWN, The Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, OX1 3QZ UK; KENNETH M. EVENSON, NIST, 325 Broadway, Boulder, CO 80303, USA.

The fine-structure transition ^{2}Π_{1/2} ← ^{2}Π_{3/2} of the free radical FO has been detected by far-infrared Laser Magnetic Resonance. The spectrum is entirely magnetic dipole and several pure rotational transitions ranging between J = \frac{1}{2} and J = 20 \frac{1}{2} have been recorded in both perpendicular and parallel polarization. These data form the most accurate spectroscopic information available for FO. A set of 272 transitions have been analyzed using a linear least-sqares fit. The Hamiltonian, which contains rotational, centrifugal distortion, spin-orbit, Λ−doubling, magnetic hyperfine and Zeeman terms, was set up in a decoupled basis set. The spin-orbit coupling constant A_{0}, the magnetic hyperfine parameters a, b_{F}, c, d and a number of g-factors have been derived experimentally for the first time. The value of A_{0} obtained from the fit is:

A_{0} = −196.108690(50) cm^{−1}.